OpenStudy (anonymous):

Help. I need the answer before the time runs out What property is shown in the equation? 0 ÷ –3 = 0 A. division into zero property B.identity property of division C.negative one property of division D. reciprocal property

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

What property is shown in the equation? 3 ∙ (6ab) = 18ab A. associative property of multiplication B. commutative property of multiplication C. identity property of multiplication D. reciprocal property of multiplication

OpenStudy (anonymous):

this is a screen shot

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

I believe that for the screenshot question, the answer is A.

OpenStudy (anonymous):

thank you. im taking another screen shot

OpenStudy (anonymous):

For This question "3 ∙ (6ab) = 18ab", i think it is B

OpenStudy (anonymous):

-72xy?

OpenStudy (anonymous):

Screenshots are the same!

OpenStudy (anonymous):

Solve. (–14.3) ∙ 6.7 ∙ (–5.8) = x ∙ (–5.8) ∙ (–14.3) A. –14.3 B. –5.8 C. 1 D. 6.7

OpenStudy (anonymous):

D -6.7

OpenStudy (anonymous):

Multiply and then move numbers to one side to find x

OpenStudy (anonymous):

Simplify: -(\frac{ 2 }{ 3 })(-9y)(\frac{ 3 }{ 2 })(8x)(1) A. 72xy B.-72xy C.72 D. 0

OpenStudy (anonymous):

For This question "3 ∙ (6ab) = 18ab", i think it is B B is commutative property of multiplication

OpenStudy (anonymous):

Okay, what about the one \[-(\frac{ 2 }{ 3 })(-9y)(\frac{ 3 }{ 2 })(8x)(0)\]

OpenStudy (anonymous):

72xy

OpenStudy (anonymous):

\[-(\frac{ 2 }{ 3 })(-9y)(\frac{ 3 }{ 2 })(8x)(1)\]

OpenStudy (anonymous):

Solve the word problem. The formula t =\[\frac{ d }{ r }\] gives the time t for a distance d traveled at a rate r. A train traveled 225 miles at a rate of 45 mi/hr. How long did it take the train to travel this far? A. 4.5 hr B. 5 hr C. 5.5 hr D. 6 hr

OpenStudy (anonymous):

This is pretty easy and I think you can do this...

OpenStudy (anonymous):

Put the numbers into the formula and do the calculation...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Solve for t. d = rt A. t = dr B. t = \[\frac{ r }{ d }\] C. t = \[\frac{ d }{r }\] D. t = d – r

OpenStudy (anonymous):

2. Solve for k. m = \[\frac{ k }{ s }\] A. k = \[\frac{ m }{ s }\] B. k = ms C. k = \[\frac{ s }{ m }\] D. k = m + s

OpenStudy (anonymous):

3. Solve for t. P = irt A. t = \[\frac{ p }{ ir }\] B. t = \[\frac{ ir }{ p }\] C. t = Pir D. t = P – i – r

OpenStudy (anonymous):

4. Solve for v. M= \[\frac{ bv }{ k }\] A. v = \[\frac{ mb }{ k }\] B. v = Mkb C. v = \[\frac{ b }{ mk }\] D. v = \[\frac{ mk }{b }\]

OpenStudy (anonymous):

5. Solve for h. a = ch + f A. h = \[\frac{ a+f }{ c }\] B. h = \[\frac{ a-f }{ c}\] C. h = a/c – f D. h = \[\frac{ a-c }{ f }\]

OpenStudy (anonymous):

that's all for right now

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

anyone?